three letters are written to different persons  what is the probability three letters go into right address in an envelope
?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Nitish
 
Total number of ways in which we can put the letters in the envelopes= 3!=6; Now, No. of letters put correctly and number of ways are as follows: 0(none) letters kept correctly= x(let) ways ; 1 letter kept correctly= 3 ways; 2 letters kept correctly=0 ways; 3 letters kept correctly= 1 way; solving further, x+3+1=6 x=2 therefore, probability that none of the letters is placed into the right envelope=2/6=1/3